# Equilibrium Matching: Generative Modeling with Implicit Energy-Based Models

### v2

@@ 2 Preliminaries: Flow Matching
Diffusion and Flow Matching models learn non-equilibrium dynamics. Flow Matching (FM), for example, learns to match the conditional velocity along a linear path connecting noise and image samples. During sampling, Flow Matching starts from pure Gaussian noise and iteratively denoises the current sample using the velocity predicted by $f$. This process is governed by a differential equation framework, in which the predicted velocity is treated as the time derivative of the desired sampling path and integrated over a total length of $1$. Formally, let $f$ denote the generative model, let $\epsilon$ be Gaussian noise, let $x$ be a real image sample from the training dataset, and let $t$ be a timestep sampled uniformly between 0 and $1$. The training objective of FM can be written as:
@@ 3.1 Training
To train an Equilibrium Matching model, we aim to construct an energy landscape in which the target gradient at ground-truth samples is zero. To do so, we first define a corruption scheme that provides a transition between data and noise. Let $\gamma$ be an interpolation factor sampled uniformly between 0 and $1$, let $\epsilon$ be Gaussian noise, and let $x$ be a sample from the training set. Denote $x_{\gamma}=\gamma x+(1-\gamma)\epsilon$ as an intermediate corrupted sample. Unlike $t$ in FM, our $\gamma$ is implicit and not seen by the model. Our goal is to define a target gradient at these intermediate samples $x_{\gamma}$ that matches an implicit energy landscape. Using a gradient direction that descends from noise to data, we write the Equilibrium Matching training objective as:
@@
Linear Decay. A natural choice for $c$ is a linear function that decays from $1$ to 0. In the energy landscape, this is equivalent to assigning noise a high gradient and making the gradient decay linearly to 0 toward the ground-truth image:
@@
Truncated Decay. Beyond linear decay, we may want the gradient to remain constant when far away from data. This leads to a truncated decay, where the target gradient stays at $1$ when $\gamma\leq a$ ($a\in[0,1)$) and then decays linearly to 0 when approaching the data manifold:
@@ 3.4 Implementation
The sampling function takes in model $f$, initial state $st$, step size $\eta$, NAG factor $\mu$, and threshold $g$, and returns the sample.
@@
while norm(grad) < g:
@@ 4 Analysis
Combining Statement 1 and Statement 2, Equilibrium Matching learns ground-truth images as local minima during training. Next, we show that sampling on an Equilibrium Matching model with gradient descent converges at a rate of $O(\frac{1}{N})$, where $N$ is the total number of sampling steps.
@@ 5.1 Image Generation
Scalability. To assess scalability, we vary training length, model size, and patch size. We report the evolution of FID over training using the XL/2 training curves. For model-size scaling, we fix the patch size to $2$ and the number of training epochs to $80$, and evaluate all four model sizes: S, B, L, and XL. For patch-size scaling, we fix the model size to B and training to $80$ epochs, and evaluate patch sizes $8$, 44, and $2$. As shown in Fig.˜3, Equilibrium Matching scales well along all axes and consistently outperforms Flow Matching under all tested configurations. These results suggest that Equilibrium Matching has strong scaling potential and is a promising alternative to Flow Matching.
@@ 5.3 Ablation Study
Hyperparameter Choices. To determine the best target landscape, we search over choices and hyperparameters for $c(\gamma)$. We use our EqM-B/2 model and 80 epochs of training for all hyperparameter experiments. We tune the step size of GD sampler and report the best result for each setting in Table˜3. The best-performing gradient landscape is truncated decay with $a=0.8$. Our results suggest that it is helpful to keep a constant target gradient at the start of the trajectory before decaying to 0. We then sweep the gradient multiplier $\lambda$ on this best-performing gradient field. As shown in Fig.˜9, a multiplier of $4$ significantly improves FID. We use this setting ($c_{\text{trunc}}$ with $a=0.8$ and $\lambda=4$) as default in our experiments.
@@ D.1 ODE Sampling as a Special Case of Optimization
which is exactly a gradient-descent update with step size $\eta=h=\tfrac{1}{N}$. Hence, with $N$ steps on a unit time horizon, the gradient-descent sampler with $\eta=\tfrac{1}{N}$ coincides with the explicit Euler ODE sampler.

### v3

@@ 2 Preliminaries: Flow Matching
Diffusion and Flow Matching models learn non-equilibrium dynamics. Flow Matching (FM), for example, learns to match the conditional velocity along a linear path connecting noise and image samples. During sampling, Flow Matching starts from pure Gaussian noise and iteratively denoises the current sample using the velocity predicted by $f$. This process is governed by a differential equation framework, in which the predicted velocity is treated as the time derivative of the desired sampling path and integrated over a total length of 11. Formally, let $f$ denote the generative model, let $\epsilon$ be Gaussian noise, let $x$ be a real image sample from the training dataset, and let $t$ be a timestep sampled uniformly between 0 and 11. The training objective of FM can be written as:
@@ 3.1 Training
To train an Equilibrium Matching model, we aim to construct an energy landscape in which the target gradient at ground-truth samples is zero. To do so, we first define a corruption scheme that provides a transition between data and noise. Let $\gamma$ be an interpolation factor sampled uniformly between 0 and 11, let $\epsilon$ be Gaussian noise, and let $x$ be a sample from the training set. Denote $x_{\gamma}=\gamma x+(1-\gamma)\epsilon$ as an intermediate corrupted sample. Unlike $t$ in FM, our $\gamma$ is implicit and not seen by the model. Our goal is to define a target gradient at these intermediate samples $x_{\gamma}$ that matches an implicit energy landscape. Using a gradient direction that descends from noise to data, we write the Equilibrium Matching training objective as:
@@
Linear Decay. A natural choice for $c$ is a linear function that decays from 11 to 0. In the energy landscape, this is equivalent to assigning noise a high gradient and making the gradient decay linearly to 0 toward the ground-truth image:
@@
Truncated Decay. Beyond linear decay, we may want the gradient to remain constant when far away from data. This leads to a truncated decay, where the target gradient stays at 11 when $\gamma\leq a$ ($a\in[0,1)$) and then decays linearly to 0 when approaching the data manifold:
@@ 3.4 Implementation
The sampling function takes in model $f$, initial state stst, step size $\eta$, NAG factor $\mu$, and threshold $g$, and returns the sample.
@@
while norm(grad) > g:
@@ 4 Analysis
Combining Statement 1 and Statement 2, Equilibrium Matching learns ground-truth images as local minima during training. Next, we show that sampling on an Equilibrium Matching model with gradient descent converges at a rate of $O(\frac{1}{N})$, where NN is the total number of sampling steps.
@@ 5.1 Image Generation
Scalability. To assess scalability, we vary training length, model size, and patch size. We report the evolution of FID over training using the XL/2 training curves. For model-size scaling, we fix the patch size to 22 and the number of training epochs to 8080, and evaluate all four model sizes: S, B, L, and XL. For patch-size scaling, we fix the model size to B and training to 8080 epochs, and evaluate patch sizes 88, 44, and 22. As shown in Fig.˜3, Equilibrium Matching scales well along all axes and consistently outperforms Flow Matching under all tested configurations. These results suggest that Equilibrium Matching has strong scaling potential and is a promising alternative to Flow Matching.
@@ 5.3 Ablation Study
Hyperparameter Choices. To determine the best target landscape, we search over choices and hyperparameters for $c(\gamma)$. We use our EqM-B/2 model and 80 epochs of training for all hyperparameter experiments. We tune the step size of GD sampler and report the best result for each setting in Table˜3. The best-performing gradient landscape is truncated decay with $a=0.8$. Our results suggest that it is helpful to keep a constant target gradient at the start of the trajectory before decaying to 0. We then sweep the gradient multiplier $\lambda$ on this best-performing gradient field. As shown in Fig.˜9, a multiplier of 44 significantly improves FID. We use this setting ($c_{\text{trunc}}$ with $a=0.8$ and $\lambda=4$) as default in our experiments.
@@ D.1 ODE Sampling as a Special Case of Optimization
which is exactly a gradient-descent update with step size $\eta=h=\tfrac{1}{N}$. Hence, with NN steps on a unit time horizon, the gradient-descent sampler with $\eta=\tfrac{1}{N}$ coincides with the explicit Euler ODE sampler.